# SMReferences for Appendix

### v3

@@ Abstract
Treatment effect estimation under unconfoundedness is a fundamental task in causal inference. In response to the challenge of analyzing high-dimensional datasets collected in substantive fields such as epidemiology, genetics, economics, and social sciences, many methods for treatment effect estimation with high-dimensional nuisance parameters (the outcome regression and the propensity score) have been developed in recent years. However, it is still unclear what is the necessary and sufficient sparsity condition on the nuisance parameters for the treatment effect to be $\sqrt{n}$-estimable. In this paper, we propose a new Double-Calibration strategy that corrects the estimation bias of the nuisance parameter estimates computed by regularized high-dimensional techniques and demonstrate that the corresponding Doubly-Calibrated estimator achieves $\sqrt{n}$-rate as long as one of the nuisance parameters is sparse with sparsity below $\sqrt{n}/\log p$, where $p$ denotes the ambient dimension of the covariates, whereas the other nuisance parameter can be arbitrarily complex and completely misspecified. The Double-Calibration strategy can also be applied to settings other than treatment effect estimation, e.g. regression coefficient estimation in the presence of diverging number of controls in a semiparametric partially linear model.
@@ 1 Introduction
This article concerns the problem of efficient estimation of a parameter of scientific interest, denoted as $\tau\equiv\tau(\theta)$ under the semiparametric framework, where $\theta$ denotes the (potentially) high-dimensional nuisance parameters. A typical example of such parameters of interest is $\tau\equiv\mathbb{E}[Y(t)]$, which denotes the mean of a potential outcome $Y(t)$ with the binary treatment $T$ set to $t\in\{0,1\}$ and is identifiable from the observed data under no unmeasured confounding (Imbens and Rubin, 2015; Hernán and Robins, 2023). Since the average treatment effect (ATE) is just the contrast $\mathbb{E}[Y(1)]-\mathbb{E}[Y(0)]$, to simplify our exposition, we refer to $\tau=\mathbb{E}[Y(1)]$ as ATE and take $t=1$ throughout. To estimate ATE under the no unmeasured confounding assumption, we need to account for the nuisance parameters $\theta$, which can be decomposed into two components $\theta\equiv(\pi,r)$: the Propensity Score (PS), denoted as $\pi(x)\coloneqq\mathbb{E}[T|X=x]$, and the Outcome Regression (OR) model, denoted as $r(x)\coloneqq\mathbb{E}[Y|X=x,T=1]$. Here $X$ denotes the $p$-dimensional covariates. Although parameters other than ATE will also be covered in this paper, we mainly focus on ATE in the Introduction.
It is now well-established in modern semiparametric theory that the (nonparametric first-order) influence function (Robins et al., 1994) of $\tau$ is of the form
@@
which gives rise to the celebrated Augmented Inverse Probability Weighting (AIPW) estimator (Robins et al., 1994; Hahn, 1998). The AIPW estimator is Doubly-Robust (DR) (Scharfstein et al., 1999; Robins and Rotnitzky, 2001) and is $\sqrt{n}$-consistent and asymptotically normal when $\pi$ and $r$ are estimated at sufficiently fast rates. To establish these convergence rates, it is essential to impose certain complexity-reducing assumptions (Robins and Ritov, 1997; Ritov et al., 2014; Liu et al., 2020, 2023) on the nuisance parameters $\pi$ and $r$.
@@
During the past decade, in response to the challenge of high-dimensional datasets with (ambient) dimension $p$ potentially much greater than the sample size $N$, a large body of literature in statistics and econometrics (Farrell (2015); Shortreed and Ertefaie (2017); van der Laan (2017); Chernozhukov et al. (2018); Ju et al. (2020); Chernozhukov et al. (2022); Avagyan and Vansteelandt (2021); Ning et al. (2020); Tan (2020b, a); Tang et al. (2022); Athey et al. (2018); Hirshberg and Wager (2021); Bradic et al. (2019b, a); Wang and Shah (2020); Dukes and Vansteelandt (2021); Sun and Tan (2022) and references therein) has been devoted to solving this problem by positing $\pi$ and/or $r$ to be sparse Generalized Linear Models (GLMs): $\pi(x)\equiv\phi(\gamma^{\top}x)$, $r(x)\equiv\psi(\beta^{\top}x)$ where $\gamma$ and/or $\beta$ has only a few non-zero coordinates, and $\phi$ and $\psi$ are two (possibly nonlinear) link functions. We further denote $s_{\pi}\coloneqq\|\gamma\|_{0}$ and $s_{r}\coloneqq\|\beta\|_{0}$ as the sparsities of $\pi$ and $r$.
@@
A natural solution in this context is to first estimate $\gamma$ and $\beta$ by high-dimensional regularized regression approaches, and then estimate $\tau$ by taking the sample average of $H(\theta)$, with $\theta$ replaced by the above estimates of $\gamma$ and $\beta$. This route was indeed taken by some of the earlier works in this direction, such as Farrell (2015) and Chernozhukov et al. (2018) [^2] Chernozhukov et al. (2018) alternatively interpreted this construction as Neyman-Orthogonalization, a terminology also adopted in later works such as Mackey et al. (2018) and Foster and Syrgkanis (2019).. In particular, the seminal Double Machine Learning (DML) estimator proposed in Chernozhukov et al. (2018) splits the entire sample into two non-overlapping subsets, one used to estimate $\gamma,\beta$ by $\ell_{1}$-regularized regression, denoted as $\hat{\gamma},\hat{\beta}$, and the other used to estimate $\tau$ by taking the sample average of $H(\hat{\theta})$, denoted as $\hat{\tau}_{\textrm{DML}}$:
@@
$\hat{\tau}_{\textrm{DML}}\coloneqq\frac{1}{n}\sum_{i=1}^{n}H_{i}(\hat{\theta}),\text{ where }\hat{\theta}=(\hat{\pi},\hat{r})\text{ and }\hat{\pi}(x)=\phi(\hat{\gamma}^{\top}x),\hat{r}(x)=\psi(\hat{\beta}^{\top}x).$ (2)
@@
Sample-splitting de-correlates the randomness in the nuisance parameter estimates $\hat{\theta}$ and in $\hat{\tau}_{\textrm{DML}}$, eliminating the reliance on Donsker-type conditions on $\theta$, which is violated for high-dimensional sparse GLMs (Chernozhukov et al., 2018). The information loss due to sample-splitting can then be restored by cross-fitting (Schick, 1986; Ayyagari, 2010; Zheng and van der Laan, 2011; Robins et al., 2013). In this paper, we will also adopt the sample-splitting strategy, and estimate $\gamma$ and $\beta$ from a separate training sample (denoted as $\mathcal{D}_{\mathrm{tr}}$, see later Section 2.1) using $\ell_{1}$-regularized generalized regression techniques. To simplify exposition, however, we omit the cross-fitting step. We refer readers to Chernozhukov et al. (2018); Jiang et al. (2022) for a more thorough discussion on the potential benefits of cross-fitting.
@@
Since one typically estimates $\pi$ and $r$ at rates $\sqrt{s_{\pi}\log p/n}$ and $\sqrt{s_{r}\log p/n}$ (Bühlmann and van de Geer, 2011), it is straightforward to show that the DML estimator $\hat{\tau}_{\textrm{DML}}$ is $\sqrt{n}$-consistent and asymptotically normal as long as [^3] Throughout this paper, we adopt the common asymptotic notation such as ,,,$\ll$,$\gg$,$\lesssim$,$\gtrsim$,$\asymp$ , ≫ , ≲ , ≳ , ≍, $o(\cdot),O(\cdot),o_{\mathbb{p}}(\cdot),O_{\mathbb{p}}(\cdot)$.
@@
This condition on $s_{\pi}\cdot s_{r}$ could be quite stringent: e.g. if $s_{\pi}\asymp\sqrt{n}/\log p$, the (DML sparsity) condition implies $s_{r}\ll\sqrt{n}/\log p$. As a result, statisticians and econometricians have been puzzled for several years by the following question:
@@
What is the sufficient and necessary, and hence minimal, sparsity condition on $\pi$ and $r$ for the parameter of interest, $\tau$, to be $\sqrt{n}$-estimable; and can we construct a $\sqrt{n}$-consistent estimator of $\tau$ under this minimal sparsity condition?
Literature overview and our contribution
Since Chernozhukov et al. (2018), to address Problem*, various attempts have been made to weaken the (DML sparsity) condition. Chernozhukov et al. (2022) showed that $s_{\pi}\wedge s_{r}\ll\sqrt{n}/\log p$ is sufficient for the DML estimator $\hat{\tau}_{\textrm{DML}}$ to be $\sqrt{n}$-consistent, yet with an additional (strong) assumption that $\|\gamma\|_{1}$ and $\|\beta\|_{1}$ are bounded,imposing (potentially unnecessary) constraints on the magnitudes of non-zero coordinates of $\gamma$ and $\beta$. Without the bounded $\ell_{1}$-norm condition, by leveraging the idea of covariate balancing (Imai and Ratkovic, 2014; Zubizarreta, 2015; Chan et al., 2016; Athey et al., 2018; Ben-Michael et al., 2021; Bruns-Smith et al., 2023) and certain structures of the linear OR and logistic-linear PS models, Bradic et al. (2019b) relaxed the (DML sparsity) condition to $s_{\pi}\ll n/\log p,s_{r}\ll\sqrt{n}/\log p$ or $s_{\pi}\ll\sqrt{n}/\log p,s_{r}\ll n^{3/4}/\log p$. Similar result to that of Chernozhukov et al. (2018) was also obtained in Smucler et al. (2019) (the rate-double-robustness part) [^4] It is worth noting that Chernozhukov et al. (2018), Smucler et al. (2019) and Bradic et al. (2019a) also considered the slightly more general approximately sparse GLMs. In order not to confuse readers, in the Introduction, we will not explicitly distinguish the approximately sparse GLMs from the exactly sparse GLMs and paraphrase the established results in the literature under exactly sparse GLMs.. But the result of Smucler et al. (2019) is not only derived for ATE, but also for a large class of parameters – often called “DR functionals” (Liu et al., 2023) – with influence functions resembling the form in (1) (Rotnitzky et al., 2021).
@@
As an extension of (Athey et al., 2018), Bradic et al. (2019a) constructed an $\sqrt{n}$-consistent ATE estimator that additionally accommodates the setting where both the OR model and the best linear projection of the PS model is sparse. It shall be noted that they also exhibited two separate $\sqrt{n}$-consistent and asymptotically normal estimators for a different parameter of interest: the “regression coefficient”, which we will discuss further in Remark 7 of Section 4.1. Bradic et al. (2019a) also established necessary conditions for $\sqrt{n}$-estimability of several parameters including ATE, which is believed to be tight. The guessed minimal sparsity condition is also inspired by this result. The difference is that the necessary condition of Bradic et al. (2019a) is stated in terms of how close the best linear approximation with roughly $\sqrt{n}$ many covariates approximates at least one of the two nuisance parameters.
@@
Tan (2020a); Ning et al. (2020); Avagyan and Vansteelandt (2021); Dukes and Vansteelandt (2021) and Smucler et al. (2019) (the model-double-robustness part) also tackled this problem, striving for $\sqrt{n}$-consistency and asymptotic normal estimators under the (minimal sparsity) condition. However, these results are also not entirely satisfactory (at least theoretically) – although they allow one of two true nuisance parameters to be arbitrarily dense, the limits of their estimates by $\ell_{1}$-regularized regression, $\hat{\pi}$ and $\hat{r}$, still need to be sparse.
@@
As one can see, despite the tremendous progress made so far, a fully satisfactory solution to Problem* remains elusive. In this paper, we address Problem* by constructing novel $\sqrt{n}$-consistent and asymptotically normal estimators, under the (minimal sparsity) condition. This new methodology is based on a Double-Calibration (DCal) strategy to correct the bias of the DML estimator $\hat{\tau}_{\textrm{DML}}$. The basic idea of DCal is also rooted in covariate balancing, which will be elucidated in detail in Section 2. It is noteworthy that the idea of using calibration to correct for the estimation bias of PS and OR has also been used in the regularized calibrated estimator of Tan (2020a). The difference is that Tan (2020a) estimated the nuisance parameters and calibrated the estimation bias simultaneously, but our approach first proposes initial estimates of PS and OR, and then calibrates these estimates via solving some optimization problems. Our new DCal estimators do not need to know which one of the two nuisance parameters, $\pi$ or $r$, is a GLM with sparsity below $\sqrt{n}/\log p$. Furthermore, the DCal estimators are not only applicable to the ATE, but also work for the regression coefficient, answering an open question raised in Bradic et al. (2019a). To the best of our knowledge, for both ATE and regression coefficient, no alternative estimators have been constructed before, that are $\sqrt{n}$-consistent and asymptotically normal under the (minimal sparsity) condition. Table 1 directly compares some of the aforementioned results with ours. However, we want to point out an important caveat of our theoretical results – we only establish $\sqrt{n}$-consistency and asymptotic normality of the DCal estimator under the (minimal sparsity) condition, but not its semiparametric efficiency (Newey, 1990; Janková and van de Geer, 2018). This remains an important question for future work.
Notation and organization of the paper

### v4

# Root-n consistent semiparametric learning with high-dimensional nuisance parameters under minimal sparsity

@@ Abstract
Treatment effect estimation under unconfoundedness is a fundamental task in causal inference. In response to the challenge of analyzing high-dimensional datasets collected in substantive fields such as epidemiology, genetics, economics, and social sciences, various methods for treatment effect estimation with high-dimensional nuisance parameters (the outcome regression and the propensity score) have been developed in recent years. However, it is still unclear what is the necessary and sufficient sparsity condition on the nuisance parameters such that we can estimate the treatment effect at $1/\sqrt{n}$-rate. In this paper, we propose a new Double-Calibration strategy that corrects the estimation bias of the nuisance parameter estimates computed by regularized high-dimensional techniques and demonstrate that the corresponding Doubly-Calibrated estimator achieves $1/\sqrt{n}$-rate as long as one of the nuisance parameters is sparse with sparsity below $\sqrt{n}/\log p$, where $p$ denotes the ambient dimension of the covariates, whereas the other nuisance parameter can be arbitrarily complex and completely misspecified. The Double-Calibration strategy can also be applied to settings other than treatment effect estimation, e.g. regression coefficient estimation in the presence of a diverging number of controls in a semiparametric partially linear model, and local average treatment effect estimation with instrumental variables.
@@ 1 Introduction
This article concerns the problem of efficient estimation of a parameter of scientific interest, denoted as $\tau\equiv\tau(\theta)$ under the semiparametric framework, where $\theta$ denotes the (potentially) high-dimensional nuisance parameters. A typical example of such parameters of interest is $\tau\equiv\mathbb{E}[Y(t)]$, which denotes the mean of a potential outcome $Y(t)$ with the binary treatment $T$ set to $t\in\{0,1\}$ and is identifiable from the observed data under no unmeasured confounding. Since the average treatment effect (ATE) is just the contrast $\mathbb{E}[Y(1)]-\mathbb{E}[Y(0)]$, to simplify our exposition, we refer to $\tau=\mathbb{E}[Y(t)]$ as the ATE (with a slight abuse of terminology) and take $t=1$ throughout. To estimate ATE under the no unmeasured confounding assumption, we need to account for the nuisance parameters $\theta$, which can be decomposed into two components $\theta\equiv(\pi,r)$: the Propensity Score (PS), denoted as $\pi(x)\coloneqq\mathbb{E}[T|X=x]$, and the Outcome Regression (OR) model, denoted as $r(x)\coloneqq\mathbb{E}[Y|X=x,T=1]$. Here $X$ denotes the $p$-dimensional covariates. Although parameters other than ATE will also be covered in this paper, we mainly focus on ATE in the Introduction.
It is now well-established in modern semiparametric theory that the (nonparametric first-order) influence function (Robins et al., 1994; Hahn, 1998) of $\tau$ is of the form
@@
which gives rise to the celebrated Augmented Inverse Probability Weighting (AIPW) estimator. The AIPW estimator is Doubly-Robust (DR) (Scharfstein et al., 1999; Robins and Rotnitzky, 2001) and is $\sqrt{n}$-consistent and asymptotically normal when $\pi$ and $r$ are estimated at sufficiently fast rates. To establish these convergence rates, it is essential to impose certain complexity-reducing assumptions (Robins and Ritov, 1997; Ritov et al., 2014; Liu et al., 2020, 2024) on the nuisance parameters $\pi$ and $r$.
@@
During the past decade, in response to the challenge of high-dimensional datasets with (ambient) dimension $p$ potentially much greater than the sample size $N$ ($p\gg N$), a large body of literature in statistics and econometrics (Farrell (2015); Shortreed and Ertefaie (2017); van der Laan (2017); Chernozhukov et al. (2018); Ju et al. (2020); Chernozhukov et al. (2022b); Avagyan and Vansteelandt (2021); Ning et al. (2020); Tan (2020b, a); Tang et al. (2023); Athey et al. (2018); Hirshberg and Wager (2021); Bradic et al. (2019b, a); Wang and Shah (2024); Dukes and Vansteelandt (2021); Sun and Tan (2022) and references therein) has been devoted to solving this problem by positing $\pi$ and/or $r$ to be sparse Generalized Linear Models (GLMs): $\pi(x)\equiv\phi(\gamma^{\top}x)$, $r(x)\equiv\psi(\beta^{\top}x)$ where $\gamma$ and/or $\beta$ has only a few non-zero coordinates, and $\phi$ and $\psi$ are two (possibly nonlinear) link functions. We further denote $s_{\pi}\coloneqq\|\gamma\|_{0}$ and $s_{r}\coloneqq\|\beta\|_{0}$ as the sparsities of $\pi$ and $r$.
@@
A natural solution in this context is to first estimate $\gamma$ and $\beta$ by high-dimensional regularized regression approaches, and then estimate $\tau$ by taking the sample average of $H(\theta)$, with $\theta$ replaced by the above estimates of $\gamma$ and $\beta$. This route was indeed taken by some of the earlier works in this direction, such as Farrell (2015) and Chernozhukov et al. (2018) [^1] Chernozhukov et al. (2018) alternatively interpreted this construction as Neyman-Orthogonalization, a terminology also adopted in later works such as Mackey et al. (2018) and Foster and Syrgkanis (2023).. In particular, the seminal Double Machine Learning (DML) estimator proposed in Chernozhukov et al. (2018) splits the entire sample into two non-overlapping subsets, one used to estimate $\gamma,\beta$ by $\ell_{1}$-regularized regression, denoted as $\hat{\gamma},\hat{\beta}$, and the other used to estimate $\tau$ by taking the sample average of $H(\hat{\theta})$, denoted as $\hat{\tau}_{\mathrm{DML}}$:
@@
$\hat{\tau}_{\mathrm{DML}}\coloneqq\frac{1}{n}\sum_{i=1}^{n}H_{i}(\hat{\theta}),\text{ where }\hat{\theta}=(\hat{\pi},\hat{r})\text{ and }\hat{\pi}(x)=\phi(\hat{\gamma}^{\top}x),\hat{r}(x)=\psi(\hat{\beta}^{\top}x).$ (2)
@@
Sample-splitting de-correlates the randomness in the nuisance parameter estimates $\hat{\theta}$ and in $\hat{\tau}_{\mathrm{DML}}$, eliminating the reliance on Donsker-type conditions on $\theta$, which are violated by high-dimensional sparse GLMs (Chernozhukov et al., 2018). The information loss due to sample-splitting can then be restored by cross-fitting (Schick, 1986; Ayyagari, 2010; Zheng and van der Laan, 2011; Robins et al., 2013). In this paper, we will also adopt the sample-splitting strategy, and estimate $\gamma$ and $\beta$ from a separate training sample (denoted as $\mathcal{D}_{\mathrm{tr}}$, see later Section 2.1) using $\ell_{1}$-regularized GLM techniques. To simplify exposition, we omit the cross-fitting step. We refer readers to Chernozhukov et al. (2018); Jiang et al. (2022) for a more thorough discussion on the potential benefits of cross-fitting.
@@
Since one typically estimates $\pi$ and $r$ at rates $\sqrt{s_{\pi}\log p/n}$ and $\sqrt{s_{r}\log p/n}$ using $\ell_{1}$-regularized GLMs (Bühlmann and van de Geer, 2011), it is straightforward to show that the DML estimator $\hat{\tau}_{\mathrm{DML}}$ is $\sqrt{n}$-consistent and asymptotically normal as long as [^2] Throughout this paper, we adopt the common asymptotic notation such as ,,,$\ll$,$\gg$,$\lesssim$,$\gtrsim$,$\asymp$ , ≫ , ≲ , ≳ , ≍, $o(\cdot),O(\cdot),o_{\mathbb{p}}(\cdot),O_{\mathbb{p}}(\cdot)$.
@@
Such a requirement on the product of the two sparsities ($s_{\pi}\cdot s_{r}$) could be quite stringent: e.g. if $s_{\pi}\asymp\sqrt{n}/\log p$, the (DML sparsity) condition implies $s_{r}\ll\sqrt{n}/\log p$. As a result, statisticians and econometricians have been puzzled by the following question:
@@
What is the sufficient and necessary, and hence minimal, sparsity condition on $\pi$ and $r$ for the parameter of interest, $\tau$, to be estimable at $1/\sqrt{n}$-rate; and can we construct a $\sqrt{n}$-consistent estimator of $\tau$ under this minimal sparsity condition?
1.1 Literature overview and our contribution
Since Chernozhukov et al. (2018), various attempts have been made to weaken the (DML sparsity) condition. Chernozhukov et al. (2022b) showed that $s_{\pi}\wedge s_{r}\ll\sqrt{n}/\log p$ is sufficient for the DML estimator $\hat{\tau}_{\mathrm{DML}}$ to be $\sqrt{n}$-consistent, yet with an additional (strong) assumption that $\|\gamma\|_{1}$ and $\|\beta\|_{1}$ are bounded, imposing (potentially unnecessary) constraints on the magnitudes of non-zero coordinates of $\gamma$ and $\beta$. Without the bounded $\ell_{1}$-norm condition, by leveraging the idea of covariate balancing (Imai and Ratkovic, 2014; Zubizarreta, 2015; Chan et al., 2016; Athey et al., 2018; Ben-Michael et al., 2021; Bruns-Smith et al., 2023) and certain structures of the linear OR and logistic-linear PS models, Bradic et al. (2019b) relaxed the (DML sparsity) condition to $s_{\pi}\ll n/\log p,s_{r}\ll\sqrt{n}/\log p$ or $s_{\pi}\ll\sqrt{n}/\log p,s_{r}\ll n^{3/4}/\log p$. Similar result to that of Chernozhukov et al. (2018) was also obtained in Smucler et al. (2019) (the rate-double-robustness part) [^3] It is worth noting that Chernozhukov et al. (2018), Smucler et al. (2019) and Bradic et al. (2019a) also considered the slightly more general approximately sparse GLMs. In order not to confuse readers, in the Introduction, we will not explicitly distinguish the approximately sparse GLMs from the exactly sparse GLMs and paraphrase the established results in the literature under exactly sparse GLMs.. But the result of Smucler et al. (2019) is not only derived for ATE, but also for a large class of parameters – conventionally termed as “DR functionals” (Liu et al., 2024) (also see Section 5 of Chernozhukov et al. (2022a)) – with influence functions resembling the form in (1) (Rotnitzky et al., 2021).
@@ 1.1 Literature overview and our contribution
As an extension of Athey et al. (2018), Bradic et al. (2019a) constructed an $\sqrt{n}$-consistent ATE estimator that additionally accommodates the setting where both the OR model and the best linear projection of the PS model is sparse. It shall be noted that they also exhibited two separate $\sqrt{n}$-consistent and asymptotically normal estimators for a different parameter of interest: the “regression coefficient”, which we will discuss further in Remark 8 of Section 4.1. Bradic et al. (2019a) also established necessary conditions for the existence of $\sqrt{n}$-consistent estimators of several parameters including ATE, which is believed to be tight. The guessed minimal sparsity condition is also inspired by this result. The difference is that the necessary condition of Bradic et al. (2019a) is stated in terms of how close the best linear approximation with roughly $\sqrt{n}$ many covariates approximates at least one of the two nuisance parameters.
@@
Tan (2020a); Ning et al. (2020); Avagyan and Vansteelandt (2021); Dukes and Vansteelandt (2021) and Smucler et al. (2019) (the model-double-robustness part) also tackled this problem, striving for $\sqrt{n}$-consistency and asymptotic normal estimators under the (minimal sparsity) condition. However, these results are also not entirely satisfactory (theoretically) – although they allow one of two true nuisance parameters to be arbitrarily dense, the limits of their estimates by $\ell_{1}$-regularized regression, $\hat{\pi}$ and $\hat{r}$, still need to be sparse.
@@
As one can see, despite the tremendous progress made so far, a fully satisfactory solution to Problem* remains elusive. In this paper, we address Problem* by constructing novel $\sqrt{n}$-consistent and asymptotically normal estimators, under the (minimal sparsity) condition. This new methodology is based on a Double-Calibration (DCal) strategy to correct the bias of the DML estimator $\hat{\tau}_{\mathrm{DML}}$. The basic idea of DCal is also rooted in covariate balancing, which will be elucidated in detail in Section 2. It is noteworthy that the idea of using calibration to correct for the estimation bias of PS and OR has also been used in the regularized calibrated estimator of Tan (2020a). The difference is that Tan (2020a) estimated the nuisance parameters and calibrated the estimation bias simultaneously, but our approach first proposes initial estimates of PS and OR, and then calibrates these estimates sequentially via solving carefully-crafted constrained optimization problems. Our new DCal estimators do not need to know which one of the two nuisance parameters, $\pi$ or $r$, is a GLM with sparsity below $\sqrt{n}/\log p$. Furthermore, the DCal estimator is not only applicable to the ATE, but also works for the regression coefficient, answering an open question raised in Bradic et al. (2019a). To the best of our knowledge, for both ATE and regression coefficient, no alternative estimators have been constructed before, that are $\sqrt{n}$-consistent and asymptotically normal under the (minimal sparsity) condition. Table 1 directly compares some of the aforementioned results with ours. However, we want to point out an important caveat of our theoretical results – we only establish $\sqrt{n}$-consistency and asymptotic normality of the DCal estimator under the (minimal sparsity) condition, but not its semiparametric efficiency (Newey, 1990; Janková and van de Geer, 2018). Since the latter begs for establishing asymptotic lower bounds with tight constants, which is a much more involved problem, we leave this important question to future works.
1.2 Notation and organization of the paper